andy pulls a boat a distance of 14.8 meters along a dock using a rope that makes an angle of 25.0o with the horizontal. If she uses a force of 910 N, calculate work?



1
Expert's answer
2020-12-29T15:29:13-0500

The horizontal componet of the force is:


"F_h = 910N\cdot \cos 25\degree \approx 824.7N"

By definition, the work done by Andy is:


"W = F_hs"

where "s = 14.8m" is the distance. Thus, obtain:


"W = 824.7N\cdot 14.8m \approx 12206\space J"

Answer. 12206 J.